How do I use the intercept form to find the equation of the line with:
1) x-intercept= -3 and y-intercept= 2
2) x-intercept= 5 and y-intercept= 6
Thank you so much for your time and consideration!

Answer by rfer(16322)

m=2-0/0--3
m=2/3
--------------
0=2/3*-3+b
b=2
--------------
y=2/3x+2
you try the second one.